i need help solving these 2 linear equation problems. they both must be solved using elimination only.
1. x - 6y = 3 2. 6x + -4y = 12
3x+ 2y = 1 9x - 6y = 18
These 2 problems must b only solved using elimination. please i need help i tried i feel like its wrong

Answer by venugopalramana(3286)

x - 6y = 3....(1) ; 3x+ 2y = 1.....(2) Let me explain elimination to you.there are 2 unknowns. we can easily solve if there is only 1 unknown.so we try to eliminate 1 unknown. say x .
see coefficient of x in both eqns.,1 and 3. find their lcm .it is 3. so multiply eqn.1 with 3 to get 3x and eqn.2 with 1 to get the same 3x.having made coefficient of x same in the 2 eqns.,add or subtract them to eliminate x.
eqn.(1)*3....3x-18y=9......(3)
eqn.(2)*1....3x+2y=1........(4)
eqn(3)-eqn.(4)......-18y-2y=9-1...or -20y=8..or y=-8/20=-2/5.
now substitute this value of y in any eqn.say eqn.(1) to get x value
x-6*(-2/5)=3...
x=3-12/5=(15-12)/5 = 3/5.
to check your answer substitute these 2 values of x and y in the given equations and check for compliance.
Eqn.(1)....3/5 -6*(-2/5)= 3/5 +12/5 =15/5 = 3 =rhs ...ok.
Eqn.(2)....3*(3/5)+2*(-2/5)=9/5-4/5=5/5=1 = rhs ...ok
now i leave the second problem for you to do as practice

6x + -4y = 12
9x - 6y = 18